OpenStudy (anonymous):

Can someone explain to me how to express 2 logarithms as a single logarithm. I thought I knew how but apparently I do not. For example: log84 + log816

hartnn (hartnn):

log property : \(\Large \log A+\log B = \log (AB)\)